OpenStudy (anonymous):

A box has 60 balls, of which 26 are red, 16 are blue, and 18 are yellow. Balls are extracted at random then returned to the box. What is the probability that after performing this test three times exactly two are red and one is yellow?

OpenStudy (welshfella):

a tree diagram would help with this one

OpenStudy (anonymous):

ok

OpenStudy (mathmath333):

is answer \(\dfrac{169}{1125}\)

OpenStudy (anonymous):

i dont how u got that answer im just thinking about the tree @welshfella said

OpenStudy (mathmath333):

i mean is the answer already given in book

OpenStudy (welshfella):

its hard to draw a tree diagram with 3 draws!!! lol

OpenStudy (anonymous):

perhaps u can attaches as a image file lol

OpenStudy (welshfella):

possible scenarios are R R Y R Y R Y R R probability of each of these is 26/60 * 26/60 * 18 / 60 work this out then multiply it by 3

OpenStudy (anonymous):

121168/180

OpenStudy (welshfella):

you have made a mistake somewhere

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

12168/216000

OpenStudy (anonymous):

169/3000

OpenStudy (welshfella):

26/60 * 26/60 * 18 / 60 = 13/30 * 13/30 * 3/10 = 507 / 9000

OpenStudy (welshfella):

= 169/1000 yes

OpenStudy (welshfella):

now you need to multiply that by 3 because of the 3 different arrangements of the balls)

OpenStudy (anonymous):

169/1000 or 169/3000?

OpenStudy (welshfella):

sorry its